OpenStudy (anonymous):

Please help The difference between negative three times a number and five more than twice the number

OpenStudy (anonymous):

difference between means - more than = + number = x

OpenStudy (anonymous):

-3 * x (5 * 2)? that doesnt look right.

OpenStudy (anonymous):

(-3 * x) (5 *2x) -3x - 2x = -5x - 5? idk.

OpenStudy (anonymous):

@satellite73 is this correct?

OpenStudy (anonymous):

@jim_thompson5910 ?

OpenStudy (anonymous):

@triciaal ?

OpenStudy (calculusxy):

\[\large -3n - (5 + 2n) \] is this the expression?

OpenStudy (anonymous):

well it looks right.. im not sure

OpenStudy (calculusxy):

well if this is the case then we an have an invisible one in front of the parenthesis \[\large -3n - 1(5 + 2n) = -3 - 5 -2n \]

OpenStudy (calculusxy):

now what you have to do is combine the like terms.

OpenStudy (anonymous):

-3 would be -3n right?

OpenStudy (anonymous):

if so, -3n - 2n = -5n - 5 :)

OpenStudy (calculusxy):

remember that you only have -3 - 5 -2n

OpenStudy (triciaal):